You have just visited your financial planner and decided you want to have $200,000 in your newborn daughter's college account by the time she is 18 years old. You feel you can start saving $400 a month right away for this. What rate of return will you need to reach your goal?

choices:

A. 6.91%
B. 8.37%
C. 9.78%
D. 10.91%
E. 11.97%
Answer #1

Answer - Option B

A 1 FV 2 NPER 3 PMT 4 Rate 5 Annual rate B $200,000 216 months $400 0.70% monthly 8.37% annual 6B 200000 =18*12 months 1 FV 2 NPER 3 PMT 4 Rate 5 Annual rate 400 =RATE(B2,-B3,,B1) =B4*12 monthly annual